# Heisenberg scaling based on population coding

### v1

@@ V $n$-tensor product representation of $\mathop{\rm SU}\,(t)$
When $\lambda=(\lambda_{1},\ldots,\lambda_{t})$ with $\lambda_{1}\leq\ldots\leq\lambda_{t}$ and $\sum_{j}\lambda_{j}=n$, $d_{\lambda}$ is given as [48, (4.46)]
@@
It is upper bounded by $(n+1)^{t(t-1)/2}$. Since $|Y_{t}^{n}|\leq(n+1)^{t-1}$, we have
$I(\mathop{\rm SU}\,(t))_{A^{n}}=R(\mathop{\rm SU}\,(t))_{A^{n}}=\log\Big{(}\sum_{\lambda^{\prime}\in Y_{t}^{n}}d_{\lambda^{\prime}}^{2}\Big{)}$
$\leq$ $\log(n+1)^{(t+1)t+(t-1)}=(t^{2}-1)\log(n+1).$ (32)
@@
which is shown in Section VI. Since $t^{2}-1$ is the number of parameters of $\mathop{\rm SU}\,(t)$. the leading term in (33) is twice of the case of state estimation case given in (5).
@@
$\sum_{\lambda^{\prime}\in Y_{d}^{n}}d_{\lambda^{\prime}}^{2}=\sum_{k=0}^{m}(2k+1)^{2}=\frac{(m+1)(2m+1)(2m+3)}{3}.$ (34)
@@
VI Proof of (33)
@@
$Y_{t}^{n}(a_{n}):=\Big{\{}\lambda\Big{|}\frac{n}{a_{n}t}\leq\lambda_{k+1}-\lambda_{k}\hbox{ for }k=1,\ldots,t-1\Big{\}}.$ (39)
@@
$d_{\lambda}=$ $\prod_{1\leq j<k\leq t}\frac{k-j+\lambda_{k}-\lambda_{j}}{k-j}\geq\frac{1}{t}\prod_{1\leq j<k\leq t}\frac{\lambda_{k}-\lambda_{j}}{t}$
$\geq$ $\prod_{1\leq j<k\leq t}\frac{n}{a_{n}t^{2}}=(\frac{n}{a_{n}t^{2}})^{\frac{(t-1)t}{2}}.$ (40)
@@
$\sum_{\lambda\in Y_{t}^{n}}d_{\lambda}^{2}\geq$ $\sum_{\lambda\in Y_{t}^{n}(a_{n})}d_{\lambda}^{2}\geq\sum_{\lambda\in Y_{t}^{n}(a_{n})}(\frac{n}{a_{n}t^{2}})^{(t-1)t}$
$=$ $|Y_{t}^{n}(a_{n})|(\frac{n}{a_{n}t^{2}})^{(t-1)t}\geq\frac{m_{n}^{t-1}}{(t-1)\!}(\frac{n}{a_{n}t^{2}})^{(t-1)t}.$ (42)
@@ Appendix A Derivations of (21) and (22)
Since $f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}$ is a pure state, the reference [43, (6)] shows the inequality
ϵitalic-$\epsilon\leq$ $2(M(|\psi\rangle,\epsilon)-1)^{s}\mathrm{Tr}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}^{1+s}$
$\leq$ $2M(|\psi\rangle,\epsilon)^{s}\mathrm{Tr}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}^{1+s}$ (44)
@@
Since $f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}$ is a pure state, the reference [41, (4.67)] shows the inequality
$\log(1-\epsilon)\leq$ $\max_{P}\log\mathrm{Tr}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}^{\frac{1}{1-s}}$

### v2

@@ 5 $n$-tensor product representation of $\mathop{\rm SU}\,(t)$
When $\lambda=(\lambda_{1},\ldots,\lambda_{t})$ with $\lambda_{1}\leq\ldots\leq\lambda_{t}$ and $\sum_{j}\lambda_{j}=n$, $d_{\lambda}$ is given as [53, (4.46)]
@@
It is upper bounded by $(n+1)^{t(t-1)/2}$. When the reference system $\mathbb{C}^{l}$ is sufficiently large in the sense of (12), the maximum mutual information is evaluated as
$I(\mathop{\rm SU}\,(t))_{A^{n}R}=\!R(\mathop{\rm SU}\,(t))_{A^{n}R}=\!\log\Big{(}\sum_{\lambda^{\prime}\in Y_{t}^{n}}d_{\lambda^{\prime}}^{2}\Big{)}$
$\leq$ $\log(n+1)^{(t+1)t+(t-1)}=(t^{2}-1)\log(n+1)$ (43)
because $|Y_{t}^{n}|\leq(n+1)^{t-1}$.
Due to (21), the optimal input state is
$\sum_{\lambda\in Y_{t}^{n}}d_{\lambda}\sqrt{\Big{(}\sum_{\lambda^{\prime}\in Y_{t}^{n}}d_{\lambda^{\prime}}^{2}\Big{)}^{-1}}|\phi_{\lambda}\rangle,$ (44)
Here, $|\phi_{\lambda}\rangle$ is the maximally entangled state on ${\cal H}_{\lambda}\otimes{\cal M}_{\lambda}$.
@@
which is shown in Section 6. Since $t^{2}-1$ is the number of parameters of $\mathop{\rm SU}\,(t)$. The leading term in (45) is twice of the case of state estimation case given in (5). Therefore, this behavior can be considered to achieve the Heisenberg scaling in the sense of mutual information.
@@
$\sum_{\lambda^{\prime}\in Y_{d}^{n}}d_{\lambda^{\prime}}^{2}=\sum_{k=0}^{m}(2k+1)^{2}$
$=$ $\frac{(m+1)(2m+1)(2m+3)}{3}.$ (46)
@@
Next, we focus on $M_{\epsilon}(\mathop{\rm SU}\,(t))_{AR}$. Due to (36), (43), and (45), $\log M_{\epsilon}(\mathop{\rm SU}\,(t))_{AR}$ behaves as
$\log M_{\epsilon}(\mathop{\rm SU}\,(t))_{AR}=(t^{2}-1)\log n+O(1).$ (50)
Since $\mathop{\rm SU}\,(t)$ is parametrized by $t^{2}-1$ parameters, $B(R)$ scales as $O(R^{t^{2}-1})$. Since $M_{\epsilon}(\mathop{\rm SU}\,(t))_{AR}$ scales as $O(n^{t^{2}-1})$, the upper bound of $R_{\epsilon}$ scales as $O(n^{-1})$.
6 Proof of (45)
@@
$Y_{t}^{n}(a_{n})$
::=:= $\Big{\{}\lambda\Big{|}\frac{n}{a_{n}t}\leq\lambda_{k+1}-\lambda_{k}\hbox{ for }k=1,\ldots,t-1\Big{\}}.$ (52)
@@
$d_{\lambda}=$ $\prod_{1\leq j<k\leq t}\frac{k-j+\lambda_{k}-\lambda_{j}}{k-j}$
$\geq$ $\frac{1}{t}\prod_{1\leq j<k\leq t}\frac{\lambda_{k}-\lambda_{j}}{t}$
$\geq$ $\prod_{1\leq j<k\leq t}\frac{n}{a_{n}t^{2}}=\big{(}\frac{n}{a_{n}t^{2}}\big{)}^{\frac{(t-1)t}{2}}.$ (53)
@@
$\sum_{\lambda\in Y_{t}^{n}}d_{\lambda}^{2}\geq\!\sum_{\lambda\in Y_{t}^{n}(a_{n})}d_{\lambda}^{2}\!\geq\!\sum_{\lambda\in Y_{t}^{n}(a_{n})}\big{(}\frac{n}{a_{n}t^{2}}\big{)}^{(t-1)t}$
$=$ $|Y_{t}^{n}(a_{n})|\big{(}\frac{n}{a_{n}t^{2}}\big{)}^{(t-1)t}\geq\frac{m_{n}^{t-1}}{(t-1)\!}\big{(}\frac{n}{a_{n}t^{2}}\big{)}^{(t-1)t}.$ (55)
@@ Appendix A Derivations of (32) and (33)
Since $f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}$ is a pure state, the reference [47, (6)] shows the inequality
ϵitalic-$\epsilon\leq$ $2(M(|\psi\rangle,\epsilon)-1)^{s}$
$\cdot\mathrm{Tr}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}^{1+s}$
$\leq$ $2M(|\psi\rangle,\epsilon)^{s}$
$\cdot\mathrm{Tr}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}^{1+s}$ (57)
@@
Since $f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}$ is a pure state, the reference [44, (4.67)] shows the inequality
$\log(1-\epsilon)$
$\leq$ $\max_{P}\log\mathrm{Tr}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}^{\frac{1}{1-s}}$